# To Shuffle or not to Shuffle: Auditing DP-SGD with Shuffling

### v2

@@ Abstract
The Differentially Private Stochastic Gradient Descent (DP-SGD) algorithm allows the training of machine learning (ML) models with formal Differential Privacy (DP) guarantees. Since DP-SGD processes training data in batches, it employs Poisson sub-sampling to select each batch at every step. However, it has become common practice to replace sub-sampling with shuffling owing to better compatibility and computational overhead. At the same time, we do not know how to compute tight theoretical guarantees for shuffling; thus, DP guarantees of models privately trained with shuffling are often reported as though Poisson sub-sampling was used.
This prompts the need to verify whether gaps exist between the theoretical DP guarantees reported by state-of-the-art models and their actual leakage. To do so, we introduce a novel DP auditing procedure to analyze DP-SGD with shuffling and show that DP models trained with this approach have considerably overestimated privacy guarantees (up to 4 times). In the process, we assess the impact on privacy leakage of several parameters, including batch size, privacy budget, and threat model. Finally, we study two common variations of the shuffling procedure that result in even further privacy leakage (up to 10 times). Overall, our work attests to the risk of using shuffling instead of Poisson sub-sampling vis-à-vis privacy leakage from DP-SGD.
@@ 1 Introduction
Without robust privacy protections, Machine Learning (ML) models can be vulnerable to inference attacks exposing sensitive information about the training data [51, 9, 58]. To mitigate these risks, Differentially Private Stochastic Gradient Descent (DP-SGD) [1] can be used to train models with Differential Privacy [24] (DP) guarantees, which provably bounds the leakage from a model so that no adversary can confidently learn (up to a privacy parameter $\varepsilon$) any individual-level information about the training data. DP-SGD is supported by many open-source libraries [60, 31, 7] and is deployed in state-of-the-art (SOTA) private models with performance increasingly approaching that of non-private models [19].
@@
DP-SGD with Shuffling. Since DP-SGD is computationally intensive, practitioners often attempt to optimize it. As DP-SGD processes training data in batches, the standard approach to select batches at each step is Poisson sub-sampling, which requires random access to the entire dataset and can be slow for large datasets [49]. As a result, recent work has begun to replace Poisson sub-sampling with shuffling the training data, and deterministically iterating over fixed-size batches [19, 53, 50]. An additional benefit is that modern ML pipelines (e.g., XLA compilation [47]) are optimized for fixed batch sizes, while Poisson sub-sampling produces different batch sizes, which can significantly slow down private training [13].
@@
Motivation. Correctly estimating the theoretical guarantees provided by DP-SGD with shuffling remains an active area of research [17, 16, 30]. By contrast, the privacy analysis of Poisson sub-sampling is well-studied and admits strong privacy amplification [5, 22]. As a result, it is common for models using DP-SGD to be trained using shuffling, even though DP guarantees are reported as though Poisson sub-sampling was used [49, 19, 40], which prompts the need to analyze whether this affects the actual privacy leakage. In the rest of the paper, we use DP-SGD (Shuffle) to denote the former and DP-SGD (Poisson) for the latter.
Chua et al. [16, 17] recently discuss theoretical lower bounds for a simplified version of DP-SGD. They show that shuffling’s lower bound privacy analysis, in some settings, can be significantly larger than the expected upper bounds from Poisson sub-sampling. However, their lower bound analysis does not cover the gradient computation process of DP-SGD. Also, they do not evaluate parameter settings used in SOTA models using shuffling [19, 40]. Independently, shuffling inputs has also attracted strong interest in the “shuffle model” [8], with (nearly) tight guarantees given for pure DP mechanisms that use shuffling through third-party servers [28, 55, 27, 6, 26]. That said, we focus on DP-SGD in the “central model,” where shuffling is commonly used in practice. Consequently, verifying the technical validity of the DP guarantees of private ML models trained with shuffling remains an open research problem.
@@
Figure 1: Largest gaps observed, for each dataset, between the empirical privacy leakage estimate $\varepsilon_{emp}$ and theoretical Poisson guarantees $\varepsilon$ when training by shuffling datasets using DP-SGD (in the Worst-Case threat model, 95% CI).
Roadmap. In this paper, we analyze the gap between the empirical privacy leakage of DP-SGD (Shuffle) and the theoretical guarantees provided by DP-SGD (Poisson) vis-à-vis different capabilities and variations of the shuffling procedure. We do so using DP auditing [21], a technique that is increasingly used to estimate and compare the empirical privacy leakage from DP mechanisms against the theoretical DP upper bounds [33, 32, 44, 43, 3, 11, 2].
@@
First, we define a simplified version of DP-SGD, which we denote as the “Batched Gaussian Mechanism” (BGM), and audit it using a novel method that builds on likelihood ratio functions. Our BGM variant is non-adaptive and simpler than the mechanism studied by Chua et al. [16], enabling us to audit it tightly in a principled way (see Section 3). Then, we extend our auditing procedure to DP-SGD (Shuffle) and evaluate the empirical privacy leakage under various threat models. We define increasingly powerful threat models, Natural, Partially-Informed, and Worst-Case, where the adversary can insert progressively more canary gradients and experiment with multiple datasets and common model architectures.
Although we focus on DP-SGD (Shuffle), our auditing framework is widely applicable to and can potentially be used with any sampling technique and mechanism. For instance, our auditing procedure can be used to identify variations within, and estimate privacy leakage for, implementations of shuffling itself.
@@
Main Findings. Our experiments highlight a substantial gap (up to 4×4$\times$4 ×) between the empirical privacy leakage observed from SOTA models [19, 40] and their claimed theoretical DP guarantees. In Figure 1, we summarize the largest gaps for each dataset we experiment with under the Worst-Case threat model. In the (relatively weaker) Partially-Informed model, we continue to observe DP violations from DP-SGD (Shuffle), especially for small batch sizes. For theoretical $\varepsilon=1.0$, training a LeNet model on FMNIST with batch sizes of 100 and 1,000 yields empirical privacy leakage estimates of $\varepsilon_{emp}=3.46$ and $\varepsilon_{emp}=0.96$, respectively.
@@
We also audit the privacy leakage from two variations of the shuffling procedure reported in prior work [49], i.e., partial shuffling and batch-then-shuffle. We find them in $2.6\%$ of public code repositories related to non-private ML training and show that these variations yield even larger privacy leakage (up to 10×10$\times$10 ×) compared to standard shuffling. For theoretical $\varepsilon=0.1$, the partial shuffling and batch-then-shuffle procedures result in an empirical privacy leakage of $\varepsilon_{emp}=0.29$ and $\varepsilon_{emp}=1.00$, respectively.
Overall, while our experimental evaluation highlights gaps in privacy leakage that can often only be observed for small batch sizes and relatively strong threat models, we stress that DP should hold for all hyper-parameters and against worst-case adversaries.
Summary of Contributions:
  • We present novel techniques to audit DP-SGD (Shuffle) using likelihood ratio functions, introducing a threat taxonomy specific to shuffling and adapting our auditing techniques accordingly.
@@
  • We are the first to audit DP-SGD (Shuffle) and to empirically show a substantial gap between the empirical privacy leakage observed from SOTA models [19, 40] and the theoretical DP guarantees, evaluating the impact of batch size and threat model on the empirical privacy leakage observed from DP-SGD (Shuffle).
@@
  • We investigate and find gaps in the privacy leakage from common variations to the shuffling procedure, namely, partial shuffling and batch-then-shuffle.
@@
  • Our work demonstrates the impact of shuffling and its variations on privacy leakage, calling into question the guarantees claimed by some SOTA models. This has important implications for the evaluation of models trained with DP-SGD (Shuffle) both in terms of privacy (i.e., privacy claims might be overly optimistic) and utility (i.e., hyper-parameters tuned to DP-SGD (Shuffle) may not be optimal for DP-SGD (Poisson) and vice-versa).
@@ 2.1 Differential Privacy (DP)
The definition of adjacent (aka neighboring) datasets often varies depending on the setting. Two common adjacency notions are add/remove and edit. The former corresponds to inserting/deleting a single record from the dataset (hence $|D|=|D^{\prime}|\pm 1$); the latter to replacing a single record with another ($|D|=|D^{\prime}|$). The difference in the size of the adjacent datasets limits the applicability of the add/remove adjacency in some settings (e.g., sampling w/o replacement), while guarantees under edit adjacency typically require roughly twice the amount of noise [49] since replacing a record is equivalent to first deleting then adding a record under the add/remove adjacency.
@@
$T(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\geq\Phi(\Phi^{-1}(1-\alpha)-\mu)$
where $\Phi$ is the standard normal cumulative distribution function.
@@
Privacy Loss Distribution (PLD). The PLD formalism [37] is a useful technique first proposed to derive tight theoretical guarantees for DP mechanisms. In recent work [43], PLD has also been used to tightly audit DP-SGD (Poisson) by using the given PLD to estimate the corresponding trade-off function. However, while the PLD for DP-SGD (Poisson) is known, deriving the PLD for DP-SGD (Shuffle) is still an active area of research [16, 17].
@@ 2.2 DP-SGD
Differentially Private Stochastic Gradient Descent (DP-SGD) [1] is a popular algorithm for training machine learning (ML) models with formal privacy guarantees. DP-SGD takes in input a dataset $D$ along with several hyper-parameters and proceeds iteratively, processing the dataset in batches and computing the gradients of samples one batch at a time. There are many strategies for sampling a batch from a dataset, e.g., Poisson sub-sampling, sampling without replacement, etc. Therefore, in general, DP-SGD can be defined with an abstract batch sampler $\mathcal{B}$ that takes in as input the dataset $D$ and the (expected) batch size $B$ and outputs batches sampled from the dataset. We report its pseudo-code in Algorithm 3 in Appendix A.
Poisson Sub-Sampling. The first implementations of DP-SGD used batches sampled through Poisson sub-sampling. With this approach, the batch sampler independently samples each record $(x,y)\in D$ with probability $q=B/|D|$ in each batch, where $B$ is the batch size. One major advantage of using Poisson sub-sampling is substantially reducing the amount of required noise via strong privacy amplification theorems [1, 22].
However, in practice, Poisson sub-sampling can be quite inefficient. For instance, the dataset cannot be fully loaded in memory when it is too large; thus, one needs to load a random batch in and out from the disk at every step, which is costly [49]. Furthermore, efficient modern ML pipelines (e.g., XLA compilation) require fixed-size batches to fully leverage GPU parallelization [13].
Shuffling. In practice, state-of-the-art DP-SGD implementations often rely on more computationally efficient sampling schemes, such as shuffling, while reporting DP guarantees as though Poisson sub-sampling was used [19]. In shuffling, the batch sampler randomly permutes the records first, then partitions the dataset into fixed-size batches. Since shuffling is already the standard in non-private training, this also simplifies implementations of DP-SGD by layering DP on top of existing non-private pipelines instead of redesigning and optimizing them from scratch. However, since the privacy guarantees provided by shuffling are not yet fully understood, this can create a discrepancy between the theoretical guarantees reported by state-of-the-art models and the actual privacy leakage, which we study in this work.
Zero-out Adjacency. While shuffling is more compatible with edit adjacency due to the fixed batch sizes, the Poisson scheme is more suitable for add/remove adjancency. Nevertheless, it is possible to derive guarantees for Poisson under the edit adjacency, although this is known to be very cumbersome [5]. Thus, comparing privacy guarantees between the different sub-sampling schemes is inherently complicated. Overall, zero-out adjacency is increasingly used to analyze the privacy of DP-SGD in practice [49, 14, 35].
@@ 3 DP Auditing
Privacy auditing broadly denotes the process of empirically estimating the privacy leakage from an algorithm. In DP, this typically involves experiments to estimate the empirical privacy guarantees, denoted as $\varepsilon_{emp}$, and compare them to the claimed theoretical guarantees ($\varepsilon$). For simplicity, we assume the empirical privacy guarantees are derived for the same $\delta$ as the claimed theoretical ($\varepsilon,\delta$)-DP. Therefore, we are technically comparing ($\varepsilon_{emp},\delta$) with ($\varepsilon,\delta$), but leave out $\delta$ throughout the paper to ease presentation.
If $\varepsilon_{emp}>\varepsilon$, the mechanism leaks more privacy than expected; whereas if $\varepsilon_{emp}\ll\varepsilon$, the theoretical bounds may be loose, or there may be substantial room for improvement to the privacy estimation procedure—i.e., the audit is not tight. Empirical guarantees are also useful for estimating the privacy loss in settings where tight theoretical guarantees are not currently known, e.g., when the adversary does not have access to intermediate updates [2, 3].
Implementing DP algorithms correctly is notoriously challenging [10, 34] and bugs in DP-SGD implementations have resulted in significantly degraded protections [34, 11] and realistic privacy attacks [20]. DP auditing can be used to detect these bugs [54, 43, 4] as well as to evaluate the optimality of privacy attacks [4, 44].
In this work, we present auditing techniques to empirically estimate the privacy loss of DP-SGD (Shuffle). By comparing it with the theoretical guarantees of DP-SGD (Poisson), we evaluate the impact of using shuffling instead of Poisson sub-sampling on the privacy leakage of DP-SGD. In the rest of this section, we introduce a general auditing procedure using an adversary and a distinguishing game. The attack’s success can then be converted into the lower-bound empirical privacy leakage estimate $\varepsilon_{emp}$.

### v3

# Meenatchi Sundaram Muthu Selva Annamalai1, Borja Balle2, Jamie Hayes2, Emiliano De Cristofaro3

@@ Abstract
The Differentially Private Stochastic Gradient Descent (DP-SGD) algorithm supports the training of machine learning (ML) models with formal Differential Privacy (DP) guarantees. Traditionally, DP-SGD processes training data in batches using Poisson subsampling to select each batch at every iteration. More recently, shuffling has become a common alternative due to its better compatibility and lower computational overhead. However, computing tight theoretical DP guarantees under shuffling remains an open problem. As a result, models trained with shuffling are often evaluated as if Poisson subsampling were used, which might result in incorrect privacy guarantees.
This raises a compelling research question: can we verify whether there are gaps between the theoretical DP guarantees reported by state-of-the-art models using shuffling and their actual leakage? To do so, we define novel DP-auditing procedures to analyze DP-SGD with shuffling and measure their ability to tightly estimate privacy leakage vis-à-vis batch sizes, privacy budgets, and threat models. Overall, we demonstrate that DP models trained using this approach have considerably overestimated their privacy guarantees (by up to 4 times). However, we also find that the gap between the theoretical Poisson DP guarantees and the actual privacy leakage from shuffling is not uniform across all parameter settings and threat models. Finally, we study two common variations of the shuffling procedure that result in even further privacy leakage (up to 10 times). Overall, our work highlights the risk of using shuffling instead of Poisson subsampling in the absence of rigorous analysis methods.
@@ 1 Introduction
To mitigate privacy risks [49, 9, 57], the Differentially Private Stochastic Gradient Descent (DP-SGD) [1] algorithm is increasingly being used to train models with Differential Privacy (DP) guarantees [24]. More precisely, DP provably bounds the leakage from a model so that no adversary can confidently learn (up to a privacy parameter $\varepsilon$) any individual-level information about the training data. DP-SGD is supported by many open-source libraries [58, 30, 7] and is deployed in state-of-the-art (SOTA) private models with performance increasingly approaching that of non-private models [19].
@@
Subsampling vs. Shuffling in DP-SGD. Since DP-SGD is computationally intensive, practitioners often attempt to optimize it. As DP-SGD processes training data in batches, the standard approach to select batches at each step is Poisson subsampling, which requires random access to the entire dataset and can be slow for large datasets [47]. As a result, recent work has often replaced that with shuffling the training data and deterministically iterating over fixed-size batches [19, 52, 48]. Additionally, modern ML pipelines (e.g., XLA compilation [45]) are optimized for fixed batch sizes, while Poisson subsampling produces different batch sizes, which can significantly slow down private training [13].
@@
While the privacy analysis of Poisson subsampling is well-studied and admits strong privacy amplification theorems [5, 22], correctly estimating DP-SGD’s theoretical guarantees when using shuffling remains an open problem [17, 16, 29]. As a result, it has become common to train models using DP-SGD with shuffling while reporting DP guarantees as if Poisson subsampling were used [47, 19, 39]. (In the rest of the paper, we use DP-SGD (Shuffle) to denote the former DP-SGD (Poisson) for the latter.) This prompts the need to analyze whether this discrepancy affects the actual privacy guarantees of state-of-the-art models using shuffling.
@@
Figure 1: Largest gaps observed for each dataset between the empirical privacy leakage estimate $\varepsilon_{emp}$ and theoretical Poisson guarantees $\varepsilon$ when training by shuffling datasets using DP-SGD.
DP Auditing. We build on the concept of DP auditing [21], which consists in estimating and comparing the empirical privacy leakage (denoted as $\varepsilon_{emp}$) from DP mechanisms against their theoretical DP upper bounds ($\epsilon$[32, 31, 43, 42, 3, 11, 2]. This involves running attacks against the mechanism – e.g., in the case of DP-SGD, inferring whether or not a sample was used to train the model (aka membership inference [49]), and using the adversary’s success to estimate $\varepsilon_{emp}$. Finding that $\varepsilon_{emp}>\epsilon$ indicates the presence of DP violations or bugs in the DP algorithm. For the auditing to be effective, it also needs not to be “loose” – i.e., if $\varepsilon_{emp}\ll\epsilon$, the audit may not be exploiting the maximum possible privacy leakage.
Research Problem. In this paper, we present, to the best of our knowledge, the first methodology to audit DP-SGD (Shuffle), aiming to analyze the gap between its empirical privacy leakage and the theoretical guarantees provided by DP-SGD (Poisson). Notably, auditing the former is significantly harder than the latter, and presenting suitable algorithms has remained a largely unaddressed research question (see Section 8). Unlike for DP-SGD (Poisson), tight theoretical upper bounds are not currently known for DP-SGD (Shuffle) [16, 17], and privacy attacks have not been evaluated against it. Therefore, it is unclear whether and how the privacy leakage can be effectively leveraged by the auditing adversary.
@@
Roadmap. We start by auditing a simplified version of DP-SGD, which we denote as Batched Gaussian Mechanism (BGM), using a novel method that builds on likelihood ratio functions. Our BGM variant is non-adaptive and simpler than the Adaptive Batch Linear Query (ABLQ) mechanism studied in prior work [16], enabling us to audit it tightly in a principled way (see Section 3). Then, we extend our auditing procedure to DP-SGD (Shuffle) and evaluate the empirical privacy leakage under various adversarial models.
@@
Our experiments highlight a substantial gap (up to $4\times$) between the empirical privacy leakage observed from SOTA models [19, 39] and their claimed theoretical DP guarantees. In Figure 1, we summarize the largest gaps for each dataset we experiment with – e.g., on the MNLI dataset, the state-of-the-art differentially private BERT model from [39] reports a theoretical $\varepsilon=3$, while our audit results in an empirical $\varepsilon_{emp}=12.7$ due to shuffling the dataset as opposed to using Poisson sub-sampling as in the theoretical analysis.
However, the gap is not uniform across all parameter settings and threat models. Specifically, for large batch sizes and weak threat models, we find that the gap is much smaller, implying that the relationship between the theoretical Poisson DP guarantees and the actual privacy leakage observed from shuffling is a complex one.
@@
Finally, we adapt our auditing procedures to audit two variations of the shuffling procedure first reported by Ponomareva et al. [47], namely, partial shuffling and batch-then-shuffle. We find them in $2.6\%$ of public code repositories related to non-private ML training and show that these variations yield even larger privacy leakage (up to $10\times$) compared to standard shuffling. For theoretical $\varepsilon=0.1$, the partial shuffling and batch-then-shuffle procedures result in an empirical privacy leakage of $\varepsilon_{emp}=0.29$ and $\varepsilon_{emp}=1.00$, respectively.
Contributions. Our work makes several contributions:
@@
  • We are the first to audit DP-SGD (Shuffle) and show that the empirical privacy leakage of SOTA models [19, 39] is substantially larger than the theoretical DP guarantees, evaluating the impact of the batch size and the adversarial model on the privacy leakage observed from DP-SGD (Shuffle). In the process, we present novel auditing techniques using likelihood ratio functions.
@@
  • We investigate and identify gaps in the privacy leakage from two common variants of the shuffling procedure, namely, partial shuffling and batch-then-shuffle.
  • Although we focus on DP-SGD (Shuffle), our auditing framework could be applied to any sampling technique and mechanism, e.g., to identify variations within and estimate privacy leakage for implementations of shuffling itself.
@@
  • Our work attests to the impact of shuffling (and its variants) on privacy leakage, calling into question the guarantees claimed by some SOTA models. This has important implications both in terms of privacy (i.e., privacy claims might be overly optimistic) and utility (i.e., hyperparameters tuned to DP-SGD (Shuffle) may not be optimal for DP-SGD (Poisson) and vice versa).
NB: we have made the source code of our auditing procedure publicly available to support reproducibility and encourage further work in this space. [^1] See https://github.com/spalabucr/audit-shuffle.
@@ 2.1 Differential Privacy (DP)
Two common notions of adjacent datasets are add/remove and edit. The former corresponds to inserting/deleting a single record from the dataset (hence $|D|=|D^{\prime}|\pm 1$); the latter to replacing a single record with another ($|D|=|D^{\prime}|$). Note that the difference in the size of the adjacent datasets limits the applicability of the add/remove adjacency in some settings (e.g., sampling w/o replacement), while guarantees under edit adjacency typically require roughly twice the amount of noise [47] since replacing a record is equivalent to first deleting then adding a record under the add/remove adjacency.
@@
$T(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\geq\Phi(\Phi^{-1}(1-\alpha)-\mu),$
$\Phi$ being the standard normal cumulative distribution function.
@@
Privacy Loss Distribution (PLD). The PLD formalism [36] is useful to derive tight theoretical guarantees for DP mechanisms. In recent work [42], PLD has also been used to tightly audit DP-SGD (Poisson) by using the given PLD to estimate the corresponding trade-off function. However, while the PLD for DP-SGD (Poisson) is known, deriving the PLD for DP-SGD (Shuffle) is still an active area of research [16, 17].
@@ 2.2 DP-SGD
Differentially Private Stochastic Gradient Descent (DP-SGD) [1] is a popular algorithm for training machine learning (ML) models with formal privacy guarantees. DP-SGD takes in input a dataset $D$ along with several hyperparameters and proceeds iteratively, processing the dataset in batches and computing the gradients of samples one batch at a time. There are several strategies for sampling a batch from a dataset, such as Poisson subsampling and sampling without replacement. In general, DP-SGD can be defined with an abstract batch sampler $\mathcal{B}$ that takes as input the dataset $D$ and the (expected) batch size $B$, and outputs batches sampled from the dataset. We report its pseudo-code in Algorithm 4 in Appendix A.
Poisson subsampling. The first implementations of DP-SGD used batches sampled through Poisson subsampling. With this approach, the batch sampler independently samples each record $(x,y)\in D$ with probability $q=B/|D|$ in each batch, where $B$ is the batch size. One major advantage of using Poisson subsampling is substantially reducing the amount of required noise via strong privacy amplification theorems [1, 22]. However, in practice, Poisson subsampling can be quite inefficient. For instance, the dataset cannot be fully loaded in memory when it is too large; thus, one needs to load a random batch in and out from the disk at every step, which is costly [47]. Furthermore, efficient modern ML pipelines (e.g., XLA compilation) require fixed-size batches to fully leverage GPU parallelization [13].
Shuffling. In practice, state-of-the-art DP-SGD implementations often rely on more computationally efficient sampling schemes, such as shuffling, while reporting DP guarantees as though Poisson subsampling was used [19]. In shuffling, the batch sampler randomly permutes the records first, then partitions the dataset into fixed-size batches. Since shuffling is already the standard in non-private training, this also simplifies implementations of DP-SGD by layering DP on top of existing non-private pipelines instead of redesigning and optimizing them from scratch. However, since the privacy guarantees provided by shuffling are not yet fully understood, this can create a discrepancy between the theoretical guarantees reported by state-of-the-art models and the actual privacy leakage, which we study in this work.
Zero-out Adjacency. While shuffling is more compatible with edit adjacency due to the fixed batch sizes, the Poisson scheme is more suitable for add/remove adjancency. Nevertheless, it is possible to derive guarantees for Poisson under the edit adjacency, although this is known to be very cumbersome [5]. Thus, comparing privacy guarantees between the different subsampling schemes is inherently complicated. Overall, zero-out adjacency is increasingly used to analyze the privacy of DP-SGD in practice [47, 14, 34].
@@ 3 DP Auditing
Privacy auditing broadly denotes the process of empirically estimating the privacy leakage from an algorithm. In DP, this involves running experiments to estimate the empirical privacy guarantees ($\varepsilon_{emp}$) and compare them to the theoretical guarantees ($\varepsilon$). For simplicity, we assume the empirical guarantees are derived for the same $\delta$ as the theoretical ($\varepsilon,\delta$)-DP; thus, we are technically comparing ($\varepsilon_{emp},\delta$) to ($\varepsilon,\delta$), but leave out $\delta$ to ease presentation. If one finds that $\varepsilon_{emp}>\varepsilon$, the mechanism leaks more privacy than expected. As mentioned earlier, if $\varepsilon_{emp}\ll\varepsilon$, the audit is not tight, i.e., the theoretical bounds are overly conservative or there may be substantial room for improvement to the privacy estimation procedure.
Implementing DP algorithms correctly is challenging [10, 33] and bugs in DP-SGD implementations have resulted in significantly degraded protections [33, 11] and realistic privacy attacks [20]. DP auditing can be used to detect these bugs [53, 42, 4] and/or evaluate the optimality of the attacks [4, 43]. Moreover, empirical guarantees are also useful for estimating the privacy loss in settings where tight theoretical guarantees are not currently known, e.g., when the adversary does not have access to intermediate updates [2, 3].
In the rest of this section, we introduce a general DP auditing procedure using an adversary and a distinguishing game. The attack’s success can then be converted into the lower-bound empirical privacy leakage estimate $\varepsilon_{emp}$.